OpenStudy (pphalke):

Including a 5% sales tax, an inn charges 189 per night. Find the ins nightly cost before the tax is added. What is the inn's nightly cost before tax is added? i need to know the formula i may be tested on this tomorrow

OpenStudy (kayders1997):

What that is the rate everyone needs to pay them no water what?

OpenStudy (kayders1997):

*matter

OpenStudy (sbuck98):

189*5%=9.45 189.00-9.45=179.55 Get it? If not I'll explain more

OpenStudy (kayders1997):

Huh?

OpenStudy (kayders1997):

You would be adding that value

OpenStudy (kayders1997):

Oh including okay yes she's right

OpenStudy (pphalke):

i need the formula for any problem like that, i might have that one on the test

OpenStudy (sbuck98):

You just multiply the percentage from the other number, and subtract that answer from the number

OpenStudy (kayders1997):

C=R(T)where C=cost after tax r=regular cost before tax T=tax

OpenStudy (kayders1997):

Sorry no C=cost of tax

OpenStudy (kayders1997):

Than you need to subtract that from the cost after tax so you can find the original rate

OpenStudy (sbuck98):

I mean yeah, it's easier to remember how I told you though. multiply, and subtract from the number...